Jerry invested $5000 in a savings account with simple 3 years. If the interest earned from the account is $600, find the rate.
Answer by stanbon(75887)

Jerry invested $5000 in a savings account with simple 3 years. If the interest earned from the account is $600, find the rate.
-----
I(t) = P*r*t
600 = 5000*r*3
----
200 = 5000r
r = 200/5000
r = 4/100 = 0.04 = 4%
--------------
Cheers,
Stan H.
======================